em 61 l10 Poi (11) What i e velocity of a beam or roms nhaw goe elects when movin perpendiculas an elect and to a magnetic and it are also to each other and have ma mirades radius the s T, respectively what The oit electron orbit if the is uusned
Answer #1

(a) If electron move undeflected then the force due to the electric and magnetic field should cancel each other.
Hence
qE = qVB
where E is electric field , V is velocity of beam , q is charge and B is the magnetic field
V = E/B = 7.7*103 / 7.5*10-3 = 1.0267*106 m/s
(b) when the electric field is turned off then electron will follow the ciruclar path and
centripetal force will be provided by the magentic field , hence
mv2 /r = qvB
r = mv/qB = ( 9.11*10-31*1.027*106) / (1.6*10-19*7.5*10-3) = 7.797*10-4 m